# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ 2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
Let $(q_{1},q_{2},q_{3})$ be the coordinate on $\mathbb{R}^{3}$. The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form. A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ on $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
@@
We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, recall the conormal bundle $L_{K}$ defined in Section 1. Then, the unit conormal bundle of $K$ is defined by
@@
Let us prepare several notations. We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$. Then, an orientation on $L_{K}$ is determined from the orientation on $\mathbb{R}^{3}$. We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$. As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [17, Remark 3.6].
We fix an orientation on $K$. Let $l\colon K\to\Lambda_{K}$ be a section which gives the zero framing of $K$ and $m\colon S^{1}\to\Lambda_{K}$ be the meridian of $K$. The conventions are as follows: $S^{1}$ is oriented as the boundary of the unit disk $D$, and $(\nu,dl(v_{l}),dm(v_{m}))$ is a positive basis of $T_{q}\mathbb{R}^{3}$ at $q\in l(K)\cap m(S^{1})$, where $\nu\in T_{q}(L_{K}\cap D^{*}_{1}\mathbb{R}^{3})$ points outward and $v_{l}$ and $v_{m}$ are positive vectors tangent to $K$ and $S^{1}$ respectively. Then, the homology classes $[l]$ and $[m]$ give a basis of $H_{1}(\Lambda_{K})$.
@@
Therefore, the homology class $[\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
[…]
We choose a basis $\{\widehat{l},\widehat{m},\widehat{s}\}$ of $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ such that $\partial(\widehat{l})=[l]$, $\partial(\widehat{m})=[m]$ and $\widehat{s}$ is represented by the fiber $S^{*}_{q}\mathbb{R}^{3}$ of any $q\in\mathbb{R}^{3}$. For this choice, we refer to the remark after the proof of Theorem 1.2 in [4, Section 3.7], which means that we use a knot isotopic to $K$ and transversal to the contact structure $\operatorname{Ker}(dq_{3}-q_{1}dq_{2}+q_{2}dq_{1})$ on $\mathbb{R}^{3}$.
@@ 2.5.2 Exact Lagrangian cobordism arising from clean intersection
Let us make a setup to deal with Question 1.1. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and tak $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$. Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. From [17, Lemma 2.3], we may assume that there exists $\varepsilon>0$ such that
@@
by moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$. We consider a diffeomorphism
@@
is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. The inclusion maps (1) are written as
$i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{+},(q,p)),$
$i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{-},(q,p)),$
for some $r_{+},r_{-}\in\mathbb{R}$ with $r_{+}>r_{-}$.
The Lagrangian filling $\varphi(L_{K_{0}})$ of $\Lambda_{K_{0}}$ in $T^{*}\mathbb{R}^{3}$ is obtained by gluing the Lagrangian cobordism $L_{\varphi}$ in ×$\mathbb{R}\times S^{*}\mathbb{R}^{3}$ and the Lagrangian filling $L_{K_{1}}$ in $T^{*}\mathbb{R}^{3}$ along $\Lambda_{K_{1}}$. Therefore, as a part of the Mayer-Vietoris exact sequence, we have a short exact sequence
[…]
Note that $H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\cong H_{1}(L_{K_{1}})=\mathbb{Z}\cdot[K_{1}]$ and $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\cong H_{1}(L_{K_{0}})=\mathbb{Z}\cdot[K_{0}]$. The left map is a surjection since $(i_{+})_{*}(\widehat{l})\in H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is mapped to a generator of $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))$. It follows that $H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is a free abelian group of rank $3$ with a basis $\{\widehat{l}_{0},\widehat{m}_{0},\widehat{s}_{0}\}$ defined by
$\widehat{l}_{0}$ $\coloneqq(i_{+})_{*}(\widehat{l})\text{ for }\widehat{l}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}}),$
$\widehat{m}_{0}$ $\coloneqq(i_{-})_{*}(\widehat{m})\text{ for }\widehat{m}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}}),$
$\widehat{s}_{0}$ [$\coloneqq[\{0\}\times S^{*}_{q}\mathbb{R}^{3}]\text{ for any }q\in\mathbb{R}^{3}.$
@@
The open embedding $\varphi^{-1}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ preserves the orientations of $L_{K_{1}}$ and $L_{K_{0}}$. Moreover, $\varphi_{*}([K_{0}])=a\cdot[K_{0}]$ in $H_{1}(\varphi(L_{K_{0}}))$ for $a\in\{1,-1\}$.
The proof is left to Appendix A. Note that the first assertion about orientations was shown in [17, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [17, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ with the triangle product. It should be possible to prove Proposition 2.7 in a way parallel to [17, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$. However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
@@
It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$. The circle $\varphi(K_{0})$ is oriented via $\varphi\colon K_{0}\to\varphi(K_{0})$. From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $[K_{1}]=a[\varphi(K_{0})]$. Consider the map induced by $i_{\pm}$ on the first homology groups
@@
A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that for some $b\in\mathbb{Z}$,
$\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0}.\end{array}$ (5)
By the map $H_{1}(L_{\varphi})\to H_{1}(\varphi(L_{K_{0}}))$ induced by $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})$, $l_{0}$ is mapped to $[\varphi(K_{0})]$ and $m_{0}$ is mapped to 00. Since $(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([l])=[K_{1}]=a[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$, the coefficient of $l_{0}$ in $(i_{-})_{*}([l])$ is $a$. This shows the first equation of (5) for some $b\in\mathbb{Z}$.
@@
$[\partial D]=(\left.F\right|_{L_{\varphi}})_{*}((i_{+})_{*}([m]))\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$
for $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$. We may assume that $D$ intersects $K_{1}$ transversely. Then, for sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set. Then, using $a\in\{\pm 1\}$ determined by $[K_{1}]=a[\varphi(K_{0})]$, we have
@@
For the first equality, note that the inclusion map $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\varphi(L_{K_{0}})$ preserves orientations by Proposition 2.7 when $\varphi(L_{K_{0}})$ is oriented via the map $\varphi\colon L_{K_{0}}\to\varphi(L_{K_{0}})$. Let $\Sigma$ be the closure of $F^{-1}(D\setminus\coprod_{i\in I}D^{\prime}_{i})$ in $L_{\varphi}$. Then,
@@
$L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle. It is restricted on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$, then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$. This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$. (Here, we need to care about the underlying orientations of spin manifolds. See the arguments in [17, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6. ∎
Lastly, let us introduce a spin structure $\mathfrak{s}^{0}_{K}$ on $\Lambda_{K}$ as in [7, Section 6.2]. When $T\Lambda_{K}$ is trivialized as an oriented vector bundle, then $\mathfrak{s}^{0}_{K}$ is given by the trivial $\mathrm{Spin}(2)$-bundle on $\Lambda_{K}$. It determines the orientations of moduli spaces which are used to define the differential of the DGAs in [4, Section 2.3.4] and [7, Section 3.6]. The ring isomorphism $\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})$ from (2) is given by

### v2

@@ 2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
Let $(q_{1},q_{2},q_{3})$ be the coordinate of $\mathbb{R}^{3}$. The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form. A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ of $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
@@
We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, recall its conormal bundle $L_{K}$ defined in Section 1. Then, the unit conormal bundle of $K$ is defined by
@@
Let us prepare several notations. We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$. Then, an orientation of $L_{K}$ is determined from the orientation of $\mathbb{R}^{3}$. We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$. As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [16, Remark 3.6].
We fix an orientation of $K$. Let $l\colon K\to\Lambda_{K}$ be a section which gives the zero framing of $K$ and $m\colon S^{1}\to\Lambda_{K}$ be the meridian of $K$. The conventions are as follows: $S^{1}$ is oriented as the boundary of the unit disk $D$, and $(\nu,dl(v_{l}),dm(v_{m}))$ is a positive basis of $T_{q}\mathbb{R}^{3}$ at $q\in l(K)\cap m(S^{1})$, where $\nu\in T_{q}(L_{K}\cap D^{*}_{1}\mathbb{R}^{3})$ points outward and $v_{l}$ and $v_{m}$ are positive vectors tangent to $K$ and $S^{1}$ respectively. Then, the homology classes $[l]$ and $[m]$ give a basis of $H_{1}(\Lambda_{K})$.
@@
Therefore, the homology class $\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
${H_{2}(\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{0}$${H_{2}(S^{*}\mathbb{R}^{3})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$${H_{1}(\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(S^{*}\mathbb{R}^{3})=0.}$
We choose a basis $\{\widehat{l},\widehat{m},\widehat{s}\}$ of $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ such that $\partial(\widehat{l})=[l]$, $\partial(\widehat{m})=[m]$ and $\widehat{s}$ is represented by the fiber $S^{*}_{q}\mathbb{R}^{3}$ of any $q\in\mathbb{R}^{3}$. For this choice, we refer to the remark after the proof of Theorem 1.2 in [3, Section 3.7], using a knot which is isotopic to $K$ and transverse to the contact structure $\operatorname{Ker}(dq_{3}-q_{1}dq_{2}+q_{2}dq_{1})$ on $\mathbb{R}^{3}$.
@@ 2.5.2 Exact Lagrangian cobordism arising from clean intersection
Let us make a setup to deal with Question 1.1. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and let $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$. Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. By moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$, we may assume that there exists $\varepsilon>0$ such that
@@
For the proof, see [16, Lemma 2.3]. We consider a diffeomorphism
@@
is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. We have the inclusion maps (1)
$\begin{array}[]{cc}i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi}),&i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi}).\end{array}$
The Lagrangian filling $\varphi(L_{K_{0}})$ of $\Lambda_{K_{0}}$ in $T^{*}\mathbb{R}^{3}$ is obtained by gluing the Lagrangian cobordism $L_{\varphi}$ in $\mathbb{R}\times S^{*}\mathbb{R}^{3}$ and the Lagrangian filling $L_{K_{1}}$ in $T^{*}\mathbb{R}^{3}$ along $\Lambda_{K_{1}}$. Therefore, as a part of the Mayer-Vietoris exact sequence, we have a short exact sequence
${0=H_{3}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\begin{matrix}H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\\ \oplus H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\end{matrix}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}})).}$
Note that $H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\cong H_{1}(L_{K_{1}})=\mathbb{Z}\cdot[K_{1}]$ and $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\cong H_{1}(L_{K_{0}})=\mathbb{Z}\cdot[K_{0}]$. The map $u$ is a surjection since $(i_{+})_{*}(\widehat{l})\in H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is mapped to a generator of $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))$. It follows that $H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is a free abelian group of rank $3$ with a basis $\{\widehat{l}_{0},\widehat{m}_{0},\widehat{s}_{0}\}$ defined by
$\begin{cases}\widehat{l}_{0}\coloneqq(i_{+})_{*}(\widehat{l})&\text{for }\widehat{l}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}}),\\ \widehat{m}_{0}\coloneqq(i_{-})_{*}(\widehat{m})&\text{for }\widehat{m}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}}),\\ \widehat{s}_{0}\coloneqq[\{0\}\times S^{*}_{q}\mathbb{R}^{3}]&\text{for any }q\in\mathbb{R}^{3}.\end{cases}$
@@
The open embedding $\varphi^{-1}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ preserves the orientations of $L_{K_{1}}$ and $L_{K_{0}}$. Moreover, $(\left.\varphi\right|_{L_{K_{0}}})_{*}([K_{0}])=a\cdot[K_{0}]$ in $H_{1}(\varphi(L_{K_{0}}))$ for $a\in\{1,-1\}$.
The proof is left to Appendix A. Note that the first assertion about orientations was shown in [16, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [16, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ together with the triangle product. It should be possible to prove Proposition 2.7 in a way parallel to [16, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$. However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
@@
It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$. The circle $\varphi(K_{0})$ in $\varphi(L_{K_{0}})$ is oriented via the diffeomorphism $\left.\varphi\right|_{K_{0}}\colon K_{0}\to\varphi(K_{0})$. From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$. Consider the map induced by $i_{\pm}$ on the first homology groups
@@
A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that the equations
$\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0},\end{array}$ (5)
hold for some $b\in\mathbb{Z}$.
By the map $H_{1}(L_{\varphi})\to H_{1}(\varphi(L_{K_{0}}))$ induced by $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})$, $l_{0}$ is mapped to $\varphi(K_{0})]$ and $m_{0}$ is mapped to 0. Since $(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([l])=[K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$, the coefficient of $l_{0}$ in $(i_{-})_{*}([l])$ is $a$. This shows that the first equation of (5) holds for some $b\in\mathbb{Z}$.
@@
[…]
with respect to the diffeomorphism $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$. We may assume that $D$ intersects $K_{1}$ transversely. Then, their algebraic intersection number is equal to $a\in\{\pm 1\}$. For sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set. The summand of $\partial D_{i}]\in H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$ for all $i\in I$ satisfy
@@
For the first equality, note that the inclusion map $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\varphi(L_{K_{0}})$ preserves orientations by Proposition 2.7 when $\varphi(L_{K_{0}})$ is oriented via the diffeomorphism $\left.\varphi\right|_{L_{K_{0}}}\colon L_{K_{0}}\to\varphi(L_{K_{0}})$. Let $\Sigma$ be the closure of $(\left.F\right|_{L_{\varphi}})^{-1}(D\setminus\coprod_{i\in I}D^{\prime}_{i})$ in $L_{\varphi}$. Then,
@@
$L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle. We restrict $\mathfrak{s}_{L_{K_{0}}}$ on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$. Then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$. This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$. (Here, we need to care about the underlying orientations of spin manifolds. See the arguments in [16, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6. ∎
Lastly, let us introduce a spin structure $\mathfrak{s}^{0}_{K}$ on $\Lambda_{K}$ as in [6, Section 6.2]. When $T\Lambda_{K}$ is trivialized as an oriented vector bundle, then $\mathfrak{s}^{0}_{K}$ is given by the trivial $\mathrm{Spin}(2)$-bundle on $\Lambda_{K}$. It determines the orientations of moduli spaces which are used to define the differential of the DGAs in [3, Section 2.3.4] and [6, Section 3.6]. The ring isomorphism $\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})$ from (2) is given by